# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 3 Discussions on inducing metrics on unital completely positive maps
Let $A$ be a unital $C^{*}$-algebra and $B$ be a $C^{*}$-algebra with a trace $\tau$. Let $L\colon A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ be a seminorm on $A\otimes_{\rm max}B^{\rm op}$. Define the extended metric $\Delta_{\tau,L}^{\rm max}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
$\Delta_{\tau,L}^{\rm max}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm max}(F),\omega_{\tau}^{\rm max}(G)).$ (3.6)
In the case where $B$ is unital, $\tau$ is amenable, and $L\colon A\otimes_{\rm min}B^{\rm op}\to[0,\infty]$ is a seminorm on $A\otimes_{\rm min}B^{\rm op}$, we may define $\Delta_{\tau,L}^{\rm min}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
$\Delta_{\tau,L}^{\rm min}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm min}(F),\omega_{\tau}^{\rm min}(G)).$ (3.7)
In the rest of this paper we have chosen to focus on the metrics $\Delta_{\tau,L}^{\rm max}$ and $\Delta_{\tau,L}^{\rm min}$. The reason for this is that we are able to show, that for a sequence of seminorms $\{L_{n}\}_{n\in\mathbb{N}}$ (defined on the appropriate spaces) satisfying certain mild assumptions, the induced sequences $\{\Delta_{\tau,L_{n}}^{\rm max}\}_{n\in\mathbb{N}}$ and $\{\Delta_{\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$ will be stable, cf. Theorem 5.3. In Section 6 we show that the assumptions needed for stability are satisfied for a large class of $C^{*}$-algebras when the seminorms are induced from spectral triples in an appropriate manner. Moreover, we also show that chaining for both $\Delta_{\tau,L}^{\rm max}$ and $\Delta_{\tau,L}^{\rm min}$ also follow for subsets of the set of trace channels under mild assumptions on $L$, cf. Theorems 5.10 and 5.12. Note that in a lot of circumstances it might be too strict to require that Equation 3.3 holds for all unital completely positive maps, or all trace channels. Indeed, even in the matrix algebra case, [GLN05] find it necessary to put restrictions on at least one of the maps involved in their analogue of Equation 3.3, see [GLN05, pg. 5]. We manage to prove that chaining is satisfied for a large class of unital completely positive maps between twisted group $C^{*}$-algebras when the seminorms are induced from spectral triples arising from proper length functions on the underlying group, cf. Theorem 6.9.
4 Tracial $C^{*}$-algebras and completely positive maps
@@
In this section we consider a general $C^{*}$-algebra $A$ and a $C^{*}$-algebra $B$ with a trace $\tau$. We will define a map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$, which we will use to characterize the linear maps $F\colon A\to B$ satisfying that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm max}B^{\rm op}$. We will call these maps trace channels. In the case where $B$ is unital and $\tau$ is amenable, Kirchberg’s Theorem, Theorem 2.2, will give that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm min}B^{\rm op}$ as well.
4.1 The map $\omega_{\tau}$ and trace channels

### v2

@@ 4. Pullback-induced metrics
Let $L:A\to[0,\infty]$ be a seminorm. We will say that $\{L_{n}\}_{n\in\mathbb{N}}$ is an upward directed sequence of seminorms adapted to $L$ when each $L_{n}$ is a seminorm on $M_{n}(\mathbb{C})\otimes A$ such that $L=L_{1}$ and $L_{n}=L_{n+1}\circ(\iota_{n}\otimes\operatorname{id}_{A})$ for every $n\in\mathbb{N}$.
Example 4.4.
Let $L\colon A\to[0,\infty]$ be a seminorm. Then there exist several possible choices of upward directed sequences of seminorms adapted to $L$. For instance, as is easily verified, we may let $\{L_{n}^{1}\}_{n\in\mathbb{N}}$ (resp. $\{L_{n}^{\infty}\}_{n\in\mathbb{N}}$) be the sequence given by
$L_{n}^{1}(x)=\sum_{i,j=1}^{n}L(a_{i,j})\quad\text{(resp. }L_{n}^{\infty}(x)=\sup\{L(a_{i,j}):1\leq i,j\leq n\}\text{)}$
for each $n\in\mathbb{N}$ and $x=\sum_{i,j=1}^{n}e_{i,j}^{(n)}\otimes a_{i,j}$ with $a_{i,j}\in A$ for all $i,j=1,\ldots,n$. A more exotic choice is the sequence $\{L_{n}^{S}\}_{n\in\mathbb{N}}$ given by
$L_{n}^{S}(x)=\sup\Big\{\sum_{i,j=1}^{n}\big|\varphi\big(e_{i,j}^{(n)}\big)\big|\,L(a_{i,j}):\varphi\in S(M_{n}(\mathbb{C}))\Big\}.$
We leave to the reader to check that $\{L_{n}^{S}\}_{n\in\mathbb{N}}$ is an upward directed sequence of seminorms adapted to $L$.
Theorem 4.5.
Let $L\colon A\to[0,\infty]$ be a seminorm and let $\{L_{n}\}_{n\in\mathbb{N}}$ be any upward directed sequence of seminorms adapted to $L$. For each $n\in\mathbb{N}$, let $d_{n}:=d_{L_{n}}$ be the extended metric on $\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ associated with $L_{n}$. In particular, $d_{1}=d_{L}$.
Then (4.1) holds for every $n\in\mathbb{N}$ and $F_{1},F_{2}\in\operatorname{UCP}(A,B)$. Thus, $\{d_{n}^{\,\rm stab}\}_{n\in\mathbb{N}}$ is stable and we have
$d_{L}^{\,\rm stab}(F_{1},F_{2})=\lim_{m\to\infty}d_{m}(\operatorname{id}_{m}\otimes F_{1},\operatorname{id}_{m}\otimes F_{2})$
for all $F_{1},F_{2}\in\operatorname{UCP}(A,B)$.
Proof.
To prove that (4.1) holds for $n\in\mathbb{N}$, let $F_{1},F_{2}\in\operatorname{UCP}(A,B)$ and set $F:=F_{1}-F_{2}$. For each $\varphi\in S(M_{n}(\mathbb{C})\otimes B)$, we can use the Hahn-Banach extension theorem for states to pick $\widetilde{\varphi}\in S(M_{n+1}(\mathbb{C})\otimes B)$ such that $\widetilde{\varphi}\circ(\iota_{n}\otimes\operatorname{id}_{B})=\varphi$. Using that $\{L_{n}\}_{n\in\mathbb{N}}$ is upward directed, we get
$\operatorname{mk}_{L_{n}}($ $\varphi\circ(\operatorname{id}_{n}\otimes F_{1}),\varphi\circ(\operatorname{id}_{n}\otimes F_{2}))$
$=\sup\big\{\big|(\varphi\circ(\operatorname{id}_{n}\otimes F))(x)\big|:x\in M_{n}(\mathbb{C})\otimes A,L_{n}(x)\leq 1\big\}$
$=\sup\big\{\big|(\varphi\circ(\operatorname{id}_{n}\otimes F))(x)\big|:x\in M_{n}(\mathbb{C})\otimes A,(L_{n+1}\circ(\iota_{n}\otimes\operatorname{id}_{A}))(x)\leq 1\big\}$
$=\sup\big\{\big|((\widetilde{\varphi}\circ(\iota_{n}\otimes\operatorname{id}_{B}))((\operatorname{id}_{n}\otimes F)(x))\big|:x\in M_{n}(\mathbb{C})\otimes A,L_{n+1}((\iota_{n}\otimes\operatorname{id}_{A})(x))\leq 1\big\}$
$=\sup\big\{\big|((\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F))((\iota_{n}\otimes\operatorname{id}_{A})(x))\big|:x\in M_{n}(\mathbb{C})\otimes A,L_{n+1}((\iota_{n}\otimes\operatorname{id}_{A})(x))\leq 1\big\}$
$\leq\,\sup\big\{\big|((\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F))(y)\big|:y\in M_{n+1}(\mathbb{C})\otimes A,L_{n+1}(y)\leq 1\big\}$
$=\operatorname{mk}_{L_{n+1}}(\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F_{1}),\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F_{2}))$
for every $\varphi\in S(M_{n}(\mathbb{C})\otimes B)$. Hence
$d_{n}(\operatorname{id}_{n}\otimes F_{1},\operatorname{id}_{n}\otimes F_{2})$ $=\sup\{\operatorname{mk}_{L_{n}}(\varphi\circ(\operatorname{id}_{n}\otimes F_{1}),\varphi\circ(\operatorname{id}_{n}\otimes F_{2})):\varphi\in S(M_{n}(\mathbb{C})\otimes B)\}$
$\leq\,\sup\{\operatorname{mk}_{L_{n+1}}(\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F_{1}),\widetilde{\varphi}\circ(\operatorname{id}_{n+1}\otimes F_{2})):\varphi\in S(M_{n}(\mathbb{C})\otimes B)\}$
$\leq\,\sup\{\operatorname{mk}_{L_{n+1}}(\psi\circ(\operatorname{id}_{n+1}\otimes F_{1}),\psi\circ(\operatorname{id}_{n+1}\otimes F_{2})):\psi\in S(M_{n+1}(\mathbb{C})\otimes B)\}$
$=d_{n+1}(\operatorname{id}_{n+1}\otimes F_{1},\operatorname{id}_{n+1}\otimes F_{2}),$
as we wanted to show. The conclusion follows then from proposition 4.2. ∎
This establishes the main stability result for pullback-induced extended metrics from seminorms. We now proceed to show that chaining for the same metrics is satisfied under reasonable assumptions.
Proposition 4.6.
Let $A$, $B$ and $C$ be unital $C^{*}$-algebras, $L$ be a seminorm on $A$ and $K$ be a seminorm on $B$. We denote by $d_{L}$ and $d^{\prime}_{L}$ the extended metrics on $\operatorname{UCP}(A,B)$ and $\operatorname{UCP}(A,C)$ associated with $L$, respectively, and by $d_{K}$ the one on $\operatorname{UCP}(B,C)$ associated with $K$.
Let $F,F_{1},F_{2}\in\operatorname{UCP}(A,B)$, and $G,G_{1},G_{2}\in\operatorname{UCP}(B,C)$. Then we have
(4.3) $d^{\prime}_{L}(G\circ F_{1},G\circ F_{2})\leq d_{L}(F_{1},F_{2}).$
Further, if $K\circ F\leq L$, then
(4.4) $d^{\prime}_{L}(G_{1}\circ F,G_{2}\circ F)\leq d_{K}(G_{1},G_{2}).$
Finally, if $K\circ F_{1}\leq L$ or $K\circ F_{2}\leq L$, then
(4.5) $d^{\prime}_{L}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d_{L}(F_{1},F_{2})+d_{K}(G_{1},G_{2})$
Proof.
First, we have
$d^{\prime}_{L}(G\circ F_{1},G\circ F_{2})$ $=\sup\{\operatorname{mk}_{L}(\varphi\circ G\circ F_{1},\varphi\circ G\circ F_{2}):\varphi\in S(C)\}$
$\leq\,\sup\{\operatorname{mk}_{L}(\phi\circ F_{1},\phi\circ F_{2}):\phi\in S(B)\}$
$=d_{L}(F_{1},F_{2}),$
that is, (4.3) holds.
Next, assume that $K\circ F\leq L$. This implies that if $a\in A$ and $L(a)\leq 1$, then $K(F(a))\leq 1$. Thus, for any $\varphi\in S(C)$, we get
$\operatorname{mk}_{L}(\varphi\circ G_{1}\circ F,\varphi\circ G_{2}\circ F)$ $=\sup\{\big|\varphi(G_{1}(F(a)))-\varphi(G_{2}(F(a)))\big|:a\in A,L(a)\leq 1\}$
$\leq\sup\{\big|\varphi(G_{1}(F(a)))-\varphi(G_{2}(F(a)))\big|:a\in A,K(F(a))\leq 1\}$
$\leq\sup\{\big|\varphi(G_{1}(b))-\varphi(G_{2}(b))\big|:b\in B,K(b)\leq 1\}$
$=\operatorname{mk}_{K}(\varphi\circ G_{1},\varphi\circ G_{2}).$
Taking the sup over all $\varphi\in S(C)$ gives that (4.4) holds.
Finally, assume that $K\circ F_{1}\leq L$ or $K\circ F_{2}\leq L$. By symmetry, it suffices to consider the case where $K\circ F_{2}\leq L$. Then, using the triangle inequality, (4.3) with $G=G_{1}$, and (4.4) with $F=F_{2}$, we get
$d^{\prime}_{L}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d^{\prime}_{L}(G_{1}\circ F_{1},G_{1}\circ F_{2})+d^{\prime}_{L}(G_{1}\circ F_{2},G_{2}\circ F_{2})\leq d_{L}(F_{1},F_{2})+d_{K}(G_{1},G_{2}),$
showing that (4.5) holds. ∎
As a consequence, chaining for stabilized extended metrics is satisfied in the following form.
Corollary 4.7.
Let $A$, $B$ and $C$ be unital $C^{*}$-algebras, $L$ be a seminorm on $A$ and $K$ be a seminorm on $B$. Let $\{L_{n}\}_{n\in\mathbb{N}}$ be an upward directed sequence of seminorms adapted to $L$ and $\{K_{n}\}_{n\in\mathbb{N}}$ be one adapted to $K$. Denote then by $d_{L}^{\,\rm stab}$, $d_{K}^{\,\rm stab}$ and ${d}_{L}^{\,{}^{\prime}\,\rm stab}$ the stabilized extended metrics on $\operatorname{UCP}(A,B)$, $\operatorname{UCP}(B,C)$ and $\operatorname{UCP}(A,C)$ associated with these sequences, cf. theorem 4.5.
Let $F,F_{1},F_{2}\in\operatorname{UCP}(A,B)$, and $G,G_{1},G_{2}\in\operatorname{UCP}(B,C)$. Then we have
(4.6) ${d}_{L}^{\,{}^{\prime}\,\rm stab}(G\circ F_{1},G\circ F_{2})\leq d_{L}^{\,\rm stab}(F_{1},F_{2}).$
Further, if $K_{n}\circ(\operatorname{id}_{n}\otimes F)\leq L_{n}$ for every $n\in\mathbb{N}$, then
(4.7) ${d}_{L}^{\,{}^{\prime}\,\rm stab}(G_{1}\circ F,G_{2}\circ F)\leq d_{K}^{\,\rm stab}(G_{1},G_{2}).$
Finally, if $K_{n}\circ(\operatorname{id}_{n}\otimes F_{1})\leq L_{n}$ or $K_{n}\circ(\operatorname{id}_{n}\otimes F_{2})\leq L_{n}$ for every $n\in\mathbb{N}$, then
(4.8) ${d}_{L}^{\,{}^{\prime}\,\rm stab}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d_{L}^{\,\rm stab}(F_{1},F_{2})+d_{K}^{\,\rm stab}(G_{1},G_{2})$
Proof.
These inequalities follow by applying proposition 4.6 at each level and using the limit formulas for $d_{L}^{\,\rm stab}$, $d_{K}^{\,\rm stab}$ and ${d}_{L}^{\,{}^{\prime}\,\rm stab}$ provided by theorem 4.5. ∎
Remark 4.8.
Assume that the seminorms $L$ and $K$ in the theorem above satisfy that $K\circ F\leq L$ for some $F\in\operatorname{UCP}(A,B)$, and that $L_{n}:=L^{\star}_{n}$ and $K_{n}:=K^{\star}_{n}$ for every $n\in\mathbb{N}$, where $\star\in\{1,\infty,S\}$, cf. example 4.4. Then it is a routine matter to check that $K_{n}\circ(\operatorname{id}_{n}\otimes F)\leq L_{n}$ holds for every $n\in\mathbb{N}$.
Before concluding this section with a couple of examples, let us point out that if $(A,L)$ is a compact quantum metric space, then $d_{L}$ is a bounded metric on $\operatorname{UCP}(A,B)$, that is, $\operatorname{UCP}(A,B)$ has a finite diameter with respect to $d_{L}$, the reason being that $S(A)$ is then known to have finite diameter with respect to $\operatorname{mk}_{L}$. Note that even if $d_{L}$ is a bounded metric, and the sequence $\{d_{n}\}_{n\in\mathbb{N}}$ is stable with $d_{1}=d_{L}$, where $d_{n}$ is an extended metric on $\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$, there is a priori no reason to expect $d_{n}$ (or even $\delta_{1_{n}\otimes L}$) to be a bounded metric on $\operatorname{UCP}(M_{n}\otimes A,M_{n}\otimes B)$ for $n\geq 2$. In general, it would be good to know conditions under which the extended metric $d_{L}$ associated with a seminorm $L$ on $A$ (resp. $d_{L}^{\,\rm stab}$ constructed in theorem 4.5) is a genuine metric on $\operatorname{UCP}(A,B)$.
Example 4.9.
Assume that $A$ admits an ergodic (strongly continuous) action $\alpha$ of a compact group $K$, and that $\kappa$ is a continuous length function on $K$ satisfying that $\kappa(k)=0$ only if $k=1_{K}$. As shown by Rieffel in [Rie98, Theorem 2.3], $(A,L_{\alpha,\kappa})$ is then a compact quantum metric space, where $L_{\alpha,\kappa}$ is the Lipschitz seminorm on $A$ defined by
$L_{\alpha,\kappa}(a)=\sup\big\{{\|\alpha_{k}(a)-a\|}/{\kappa(k)}:k\in K\setminus\{1_{K}\}\big\}.$
Setting
$A_{\alpha,\kappa}:=\{a\in A:L_{\alpha,\kappa}(a)\leq 1\}=\{a\in A:\|\alpha_{k}(a)-a\|\leq\kappa(k)\text{ for all }k\in K\},$
we get a bounded metric $\operatorname{mk}_{L}$ on $S(A)$, given by
$\operatorname{mk}_{L}(\varphi,\psi)=\sup\{|\varphi(a)-\psi(a)|:a\in A_{\alpha,\kappa}\},$
and an associated bounded metric $d_{L_{\alpha,\kappa}}$ on $\operatorname{UCP}(A,A)$. The sequence $\{\delta_{1_{n}\otimes L_{\alpha,\kappa}}\}_{n\in\mathbb{N}}$ is stable by proposition 4.1. Concerning the chaining property from proposition 4.6 (with $A=B=C$), it will apply whenever $F_{1}$ or $F_{2}$ in $\operatorname{UCP}(A,A)$ is an $L_{\alpha,\kappa}$-contraction. As an example of such a map, let us check that this property holds for any $F\in\operatorname{UCP}(A,A)$ such that $F\circ\alpha_{k}=\alpha_{k}\circ F$ for all $k\in K$. Indeed, for every $k\in K$ and $a\in A$, we then have
$\|\alpha_{k}(F(a))-F(a)\|=\|F(\alpha_{k}(a)-a)\|\leq\|F\|\,\|\alpha_{k}(a)-a\|=\|\alpha_{k}(a)-a\|,$
since $\|F\|=\|F(1_{A})\|=1$. Thus, for every $a\in A$, we get
$L_{\alpha,\kappa}(F(a))=\sup_{k\neq 1_{K}}\big\{{\|\alpha_{k}(F(a))-F(a)\|}/{\kappa(k)}\big\}\leq\sup_{k\neq 1_{K}}\big\{{\|\alpha_{k}(a)-a\|}/{\kappa(k)}\big\}=L_{\alpha,\kappa}(a),$
as we wanted to verify.
To be more concrete, let us assume that the compact group $K$ is abelian. Let then $G=\widehat{K}$ be the dual group of $K$, pick $\sigma\in Z^{2}(G,\mathbb{T})$, and let $\alpha$ be the ergodic dual action of $K$ on $A:=C_{r}^{*}(G,\sigma)$ [Rie04], determined for each $k\in K$ by $\alpha_{k}(\lambda_{g}^{\sigma})=\langle k,g\rangle\,\lambda_{g}^{\sigma}$ for all $g\in G$. Note that each $\alpha_{k}$ is the multiplier of $A$ associated with the positive definite function $g\mapsto\langle k,g\rangle$ on $G$. Hence, for each $\phi\in P_{1}(G)$, the associated multiplier $M_{\phi}\in\operatorname{UCP}(A,A)$ commutes with every $\alpha_{k}$, and is therefore an $L_{\alpha,\kappa}$-contraction, regardless of the choice of $\kappa$. Also, by Bochner’s Theorem, every $\phi\in P_{1}(G)$ is the Fourier-Stieltjes transform of a unique probability Borel measure on $K$.
Example 4.10.
As a special case of the previous example, let $\mathbb{Z}_{n}=\{0,1,\ldots,n-1\}$ denote the cyclic group with $n$ elements and set $\zeta:=e^{i\frac{2\pi}{n}}$. Consider the ergodic action $\alpha$ of $K=\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ on $A:=B(\ell^{2}(\mathbb{Z}_{n}))\simeq M_{n}(\mathbb{C})$ implemented by the discrete Weyl operators [Wat18, Subsection 4.1.2], i.e.,
$\alpha_{(a,b)}(X)=W_{a,b}XW_{a,b}^{*}\quad\text{for all }X\in A$
for each $(a,b)\in K$, where $W_{a,b}=U^{a}V^{b}$, $U$ and $V$ being the unitary operators on $\ell^{2}(\mathbb{Z}_{n})$ satisfying $U(e_{j})=e_{j+1}$ and $V(e_{j})=\zeta^{j}e_{j}$ for $j=0,1,\ldots,n-1$, and $\{e_{j}\}_{j=0}^{n-1}$ denoting the standard basis of $\ell^{2}(\mathbb{Z}_{n})$. The simplest choice of a length function on $K$ is given by $\kappa(a,b)=1$ if $(a,b)\neq(0,0)$, while $\kappa(0,0)=0$. The associated Lipschitz seminorm $L_{\alpha,\kappa}$ on $A$ is then given by
$L_{\alpha,\kappa}(X)=\max\{\|W_{a,b}X-XW_{a,b}\|:(a,b)\in K\}.$
Linear maps from $A$ into itself that commute with all $\alpha_{a,b}$ are called Weyl-covariant in [Wat18]. A Weyl-covariant map $F$ in $\operatorname{UCP}(A,A)$ can be equivalently described as a mixed Weyl-unitary, that is, $F$ is of the form
$F(X)=\sum_{(a,b)\in K}p(a,b)W_{a,b}XW_{a,b}^{*}$
for some probability distribution $p:K\to[0,1]$, cf. [Wat18, Corollary 4.15]. Thus we get that the chaining property for $d_{L_{\alpha,\kappa}}$ is satisfied whenever $F_{1}$ or $F_{2}$ is a mixed Weyl-unitary.
Finally, note that $K$ is its own dual and that $A$ is *-isomorphic to $C_{r}^{*}(K,\sigma)$, where $\sigma\in Z^{2}(G,\mathbb{T})$ is given by $\sigma((a,b),(a^{\prime},b^{\prime}))=\zeta^{ba^{\prime}}$. Thus we could also have reached the above conclusion by making use of the second part of example 4.9.
Example 4.11.
Let $G$ be a countable discrete group, $l\colon G\to[0,\infty)$ be a proper length function, and $\sigma\in Z^{2}(G,\mathbb{T})$, so we can form the spectral triple $(A,\ell^{2}(G),\partial_{l})$, where $A:=C_{r}^{*}(G,\sigma)$. Then we get a sequence $\{\delta_{1_{n}\otimes L_{\partial_{l}}}\}_{n\in\mathbb{N}}$ which is stable, cf. proposition 4.1. If $(A,L_{\partial_{l}})$ is a compact quantum metric space, then $d_{L}$ is a bounded metric. For examples where this is known to hold, see, for instance, [Rie02, OR05, CR17, LW21, Far+24]. Moreover, as shown in lemma 2.12, if $\phi\in P_{1}(G)$, then $M_{\phi}\in\operatorname{UCP}(A,A)$ is a $L_{\partial_{l}}$-contraction; thus this class of unital completely positive maps on $A$ can be used when applying the chaining property from proposition 4.6.
5. Tracial $C^{*}$-algebras and completely positive maps
@@
In this section we consider a general $C^{*}$-algebra $A$ and a $C^{*}$-algebra $B$ having a trace $\tau$. We will define a map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$, which is an analogue of the Choi-Jamiołkowski isomorphism when $A$ and $B$ are matrix algebras. Further, we will characterize the linear maps $F\colon A\to B$ satisfying that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm max}B^{\rm op}$ and call these maps trace channels. In the case where $B$ is unital and $\tau$ is amenable, Kirchberg’s Theorem, theorem 2.2, will give that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm min}B^{\rm op}$ as well, cf. theorem 5.15.
5.1. The map $\omega_{\tau}$ and trace channels